A1.9-MeV (kinetic energy) proton enters a 0.29-t field, in a plane perpendicular to the field. what is the radius of its path?
Answer #1

Kinetic energy

KE=(1/2)mv2

(1.9*10^6*1.6*10-19)=(1/2)*(1.67*10-27)*v2

v=1.91*107 m/s

magnetic force =centripetal force

qvB=mv2/r

=>r=mv/qB =(1.67*10-27)*(1.91*107)/(1.6*10-19)*0.29

r=0.6867 m
Answer #2

For a proton
Magnetic field B=0.29T
Kinetic energy
K=1.9 MeV= 1.9*1.6*10^-19*10^6=1.6*1.9*10^-13=3.04
Answer #3

the formula is r=\frac{mv}{qB}

                     given \frac{1}{2}mv^{2}= \frac{1}{2}mv^{2}=1.9*1000000*1.6*10^{-19}

                             here 9.11*10^{-31}

v=8.17*10^{3} m/s

there fore r=\frac{9.11*10^{-31}*8.17*10^{3}}{1.6*10^{-19}*0.29}

r=1.6*10^{11} m